Two planets A and B have radii R and R/n^2 respectively what is the ratio of the time periods of their revolution
1
Expert's answer
2020-02-18T09:43:15-0500

According to orbital law 


"\frac{T_A^2}{T_B^2}=\frac{R_A^3}{R_B^3}"

"\frac{T_A^2}{T_B^2}=\frac{n^6R^3}{R^3}=n^6"

Thus,


"\frac{T_A}{T_B}=n^3"